OpenStudy (anonymous):

The sum of two consecutive integers is 15. What are the numbers?

OpenStudy (anonymous):

7,8

OpenStudy (anonymous):

8 and 7

OpenStudy (anonymous):

8 and 7

OpenStudy (asnaseer):

give the first number a symbol, say 'x', then you can write the problem as:\[x+(x+1)=15\]so:\[2x+1=15\]\[2x=14\]\[x=7\]so answer is 7 and 8

OpenStudy (anonymous):

x+y=15 y=x+1 x+x+1=15 2x=14 x=7 y=7+1 y=8 7 and 8